LHC 2012 - latest news

Week 26: Machine coordinators: Bernhard Holzer, Jan Uythoven

Plan for the week: Technical Stop until Friday 17:00, recovery and back to physics over the weekend.

 

Date

Summary

Link to report from morning meeting

Monday 2 July 06:00h injection 480 bunches, ramp squeeze, stable beams 9:00 meeting
Sunday 1 July
  • 01:09 Totem alignment finished (fast!). Also move Alpha in.
  • 01:56 Dump for the leap second.
  • 04:18 Ramp for loss maps
  • 05:53 Off-momentum loss maps in collisions with RP in
  • 06:46 Electrical perturbation (EDF) switched OFF 2 RF modules, 5 sectors (S12, S23, S67, S78, S81) and the dipoles of ALICE and LHCb
  • 07:30 pre-cycle collimator steps to be recovered
  • 10:53 Collimator problem solved, but still power converter problem RQ4.L3
  • 17:50h loss maps in squeeze / asynch dump test
  • 23:00h injection 84 bunches stable beams
9:00 meeting
Saturday 30 June
  • 04:45 problem with RQX.L8 solved
  • 07:00 injection probe beam
  • XPOC problems
  • 10:00 Ramp with 2 probes: loosing most of the beam on the ramp
  • 12:20 Ramp with 2 probes, again losses, attempt of chroma measurement in collision
  • 14:45 Scraping nominal beam for BPMS checks: ok
  • 16:30 BPMS interlock tests with bumps: ok
  • Recuperating RF settings
  • 19:15 BPMS checks with 50 ns, 12 bunches: ok
  • 20:03 Start ramp 3 bunches for Totem alignment
  • 20:30 Dump in squeeze, electrical perturbation 400 kV line
  • 22:09 Start second ramp for Totem alignment
  • 23:10 Totem starting alignment

9:00 meeting
Friday 29 June
  • 19:20 Technical Stop finished, machine closed
  • 20:00h trip of RQX.L8 investigation / repair (signal cable problem)
8:30 meeting
Thursday 28 June
  • Technical Stop
 
Wednesday 27 June
  • Technical Stop
 
Tuesday 26 June
  • Technical Stop
 
Monday 25 June
  • Overnight: measurements of dynamic aperture and non-linearities
  • 6:30 beams dumped, machine ramped down, preparaton for technical stop.
8:30 meeting
MD summary
technical stop planning